OpenStudy (anonymous):

what are the zeros of the function f(x)=x^2+x-30 3,-10 -3,10 -5,6 5,-6

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

=x^2+6x-5x-30 x(x+6)-5(x+6) =(x+6)(x-5)